OpenStudy (anonymous):

What is a possible value for the missing term of the geometric sequence? 1,250, ___, 50,

OpenStudy (anonymous):

Use \[ \Large a_1=a_nr^{n-1}\] let \(a_1=1250\) then \[\Large a_3=a_1q^2 \\ \Large 50=1250q^2 \]

OpenStudy (anonymous):

wait im still kind of confused where do i plug in the values?

OpenStudy (anonymous):

Well, can you evaluate q from what I have given you above? Pardon me by the way, you can set q=r for this, I am european, we set q rather than r for ratio. It means the same thing in the above.

OpenStudy (anonymous):

\[\Large a_n=a_1r^{n-1} \] So, this time I wrote it better, you can see that when you choose various numbers for \(n\) you can derive all the necessary parts of the GP from the problem set.

OpenStudy (anonymous):

So I assume that \(a_1\) is equal to the first part given above \(a_1=1250\) we want to figure out what \(a_2\) is, the blank spot above. We don't know that yet, however we can already write the formula down like this at the moment \[\Large a_n = 1250r^{n-1} \] So we need to figure out what r is in order to get to \(a_2\) we do that by computing \(a_3=50\) \[\Large a_3=a_1r^2 \\ \Large 50=1250r^2 \]

OpenStudy (anonymous):

oh okay so would i divide both sides? by 50

OpenStudy (anonymous):

you want to solve for \(r\)

OpenStudy (anonymous):

wait so im trying to isolate r?

OpenStudy (anonymous):

As soon as you have \(r\) you can compute any number in the GP

OpenStudy (anonymous):

yes, exactly. For any Geometric Problem, the ratio is important, that's all what it is about. A geometric progression is defined through the ratio, so as soon as you have found that ratio you're able to make forecasts about it.

OpenStudy (anonymous):

oh okay so then how would i start off? im kind of new to this

OpenStudy (anonymous):

would i square it since it's r^2?

OpenStudy (anonymous):

\[\Large 50=1250r^2 \\ \Large r=\sqrt{\frac{50}{1250}}= \pm 0.2 \]

OpenStudy (anonymous):

ohh 0.2? would be the answer, what formula is that?

OpenStudy (anonymous):

The answer isn't given by r yet alone, you need to compute \(a_2\)

OpenStudy (anonymous):

is 0.2 r? is that just the formula for finding r?

OpenStudy (anonymous):

\[\Large a_n=1250\cdot (0.2)^{n-1} \]

OpenStudy (anonymous):

So what is \(a_2\) ?

OpenStudy (anonymous):

250?

OpenStudy (anonymous):

i did 1250* 50

OpenStudy (anonymous):

you plug in a two into the equation above, that gives you your answer.

OpenStudy (anonymous):

so how would i set that up? 150 x 2? or am i wrong

OpenStudy (anonymous):

250 yes.

OpenStudy (anonymous):

excuse me that I didn't agree before, I was distracted.

OpenStudy (anonymous):

I missed your reply above.

OpenStudy (anonymous):

oh okay thankkkk youu so much! :)

OpenStudy (anonymous):

you're welcome